OpenStudy (anonymous):

What is the equivalent resistance of a parallel circuit with a 5.0 ohms resistor, a 2.5 ohms resistor, and a 10.0 ohms resistor? A. 0.10 ohms B. 1.4 ohms C. 0.83 ohms D. 10 ohms

OpenStudy (michele_laino):

the requested resistance , is given by the subsequent formula: \[\Large \begin{gathered} {R_{EQUIVALENT}} = \frac{1}{{\frac{1}{{{R_1}}} + \frac{1}{{{R_2}}} + \frac{1}{{{R_3}}}}} = \hfill \\ \hfill \\ = \frac{1}{{\frac{1}{5} + \frac{1}{{2.5}} + \frac{1}{{10}}}} = ...ohms \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

ok! we get 1.428! so choice B is our solution?

OpenStudy (michele_laino):

yes! correct!

OpenStudy (anonymous):

yay! thanks!

OpenStudy (michele_laino):

:)